Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day12


Crew 99 Log Book for February 9, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: EVA #20, 11:00-11:45; EVA #21, 12:15-12:45; EVA #22, 14:30-16:30

Brief Narrative of Field Mission Results: EVA planner; endolith sample

collection; geological data collection

EVA Data/Interpretations: As per EVA reports

Engineering/Hab Maintenance: Nothing major to report.

Report Transmission Schedule: Engineering, EVA #20, EVA #21, EVA #22, Commander's, Cook's, Journalist's. By 8:30pm

Plans for Tomorrow: More science EVAs! Also, launching tethered balloon tomorrow evening

Inventory: None

Support Requested: None

Miscellaneous: N/a


Commander's Report

Donna Viola, Crew 99, Commander

Today, as our rotation draws ever nearer to a close, we had another busy day of EVAs. Andie wrapped up the field tests of her EVA planner in the morning, first heading out with Kevin, then later with Sukrit, as they recalculated their route after coming across a contingency - a giant boulder that blocked their path! In the afternoon, Heidi and I went out to collect endolith samples and collect data for a geology project - all while exploring a really awesome canyon!

Max spent most of the day working on plans for his balloon launch. The winds just won't cooperate, so he is planning on sending up a tethered balloon tomorrow evening. Tonight is looking like another clear evening, so the astronomers are planning on heading out into the cold to spend some time in the observatory.

The rest of the team wishes them good luck and clear skies - from the warmth of the Hab!


Engineering Report

Kevin Newman Reporting

Generator/Electricity:


· Kitty (Diesel Generator) (Used? Yes/No): Yes


· Honey (Gas Generator) (Used? Yes/No): No


· Battery Charge Level: Checked by DG


o 10.59 V everstart


o 12.37 V armor plate


o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:


· Diesel (Full/0.75/0.5/0.25/Empty): Full


· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.70


· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 15 Gallons


Notes/Comments:


Oil:


· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG


· Honey Oil Quantity (# of Quarts in storage): Leave it to DG


· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench


· Car Oil Quantity (# of Quarts in storage):


Water Status:


· Outside Potable Water Tank Level (inches from bottom): 14


· Trailer Potable Water Tank Level (inches from bottom): 5/6 Full


· Hab Potable Water Tank Level (inches from bottom): 12


· Potable Water Meter Reading: 418609


Notes/Comments:


GreenHab:


· Condition of plants in Tank 1: Nothing


· Condition of plants in Tank 2: Ok


· Condition of plants in Tank 3: Good


· Greenhab notes/ comments:


Observatory:


· Telescope Used? (Yes/no) Yes


Observatory Notes: UPS in observatory has dead backup battery. Replaced with black UPS from upstairs hab, tested and working properly on battery power.


Transportation:


Hab Car (NOT RENTAL)


· Used (yes/no): No


· Oil Checked (yes/no): No


· Tire Status: Not checked


· Odometer at end of day: 123611.6


· Notes/Comments on Hab Car:


ATV 1 (Opportunity):


· Used: Yes


· Oil Checked: No


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


ATV 2 (Spirit):


· Used: Yes


· Oil Checked: No


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


ATV 3 (Viking I):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


ATV 4 (Viking II):


· Used: No


· Oil Checked: No


· Fuel Consumed: Replaced


· Tire Status: Filled


· Comments:


Heating and Ventilation:


· Thermostat setting upstairs (Farenheit): 66


· Actual temperature upstairs (Farenheit): 68


· Thermostat setting downstairs (Farenheit): 55


· Actual temperature downstairs (Farenheit): 58


Computers/Networking Infrastructure:


Notes/comments on internet/computers?: Behaving well


General Engineering

1. Notes/Comments: N/a


Astronomy Report


Sukrit Ranjan Reporting

We had a productive evening at the observatory last night.

Drift rate measurements
On Tuesday night, Max had realigned the telescope by pointing the telescope spin axis in the direction of true north as measured by asterisms in the vicinity of Polaris. Tonight we tested that alignment. On 2/4/11 we fed 2 alignment stars and 4 calibration stars into the telescope to achieve a drift rate of .35 arcsec/second at M45. Tobay we were able to put 2 alignment stars and 2 calibration stars into the telescope before it went mad Greek (see below section "Observatory of Babel"). We achieved drift rates of .11 arcsec/second when imaging near Regulus, and drift rates of .21 arcsec/second near M45. It should be noted that the image of M45 displays tracks with a slight shift in the middle of the track, suggesting the telescope was perturbed during observations. We estimate the drift rates have decreased by a factor of 2-3 due to our alignment, but further work is necessary. Peter Detterline has suggested a procedure that we will be trying.

Webcam Functionality
We successfully connected the webcam to the telescope and imaged Saturn. A useful note for future astronomers: it appears that to focus the webcam, you need to go in the "out" direction from visual focus. We will work on post-processing tomorrow.

One minor glitch: We were able to change the settings on the webcam to achieve better image quality (mostly by increasing the FPS). After changing the settings, we hit "OK", or "Save" and then "OK". However, when we then hit "Video" to start recording, the device would revert to it's original settings. We'd go back to change the settings and find them back at their starting values. We are unsure what we are doing wrong.

Observatory of Babel
We continued to experience the problem of the hand controller mysteriously switching to Greek characters (at some points it almost seemed like it had some other kind of character, maybe Chinese?)
  • Initially when I started, it was blank. It did not turn on despite heating from my hand.
  • Later, Max brought out a space heater. After heating with the space heater, unplugging the controller and turning off the telescope, and plugging in the controller and turning on the telescope, the hand controller screen started up in English -- eventually. Initially, it was solid black, and after a while, it showed English characters -- but the first character was missing. Turning off and on and further heating restored functionality.
  • I managed to get through two alignment stars and two calibration stars like this. Then the telescope started displaying the bottom line in Greek. Then the whole screen went Greek. Despite heating this did not change. I gave up, settled on just 2 calibration stars, and just steered the telescope manually to targets
  • Later, Max found that heating combined with an operation that forced the screen to refresh was able to switch the telescope back into English.
  • A note: once you select "add a calibration star" you cannot abort. At least, we could not abort the third star that we had been working on before the Greek problem happened.
  • Peter, do you need more Greek photos? I can try to post to my MIT public directory--that might work better than Apple's MobileMe system.
Observatory cover popping out
The observatory sliding roof cover popped out when pulling it open at the beginning of last night. One of the wheel pairs was no longer locked onto the track.The fix adopted was to remove the rubber stop at the base, move the cover down the tracks, pop it back on, slide it up, and reinstall the rubber stop. This seems to have worked.

EVA 20 and 21 Report

Andie Gilkey Reporting

Time: 11:30-12; 12:30-1

Crew Members:

EVA 20: Andie Gilkey, Kevin Newman

EVA 21: Andie Gilkey, Sukrit Ranjan

Site Location:  Thielges Flats, UTM 518500 4251100

Transit Mode: foot

Objectives: Use EVA mission planner to plan mission and contingency while in the field

Results: Mission planner was successfully used to find the site and re-plan in the field

Lessons Learned:  It is difficult to see the iPad screen in the bright sunlight



EVA 22 Report

Heidi Beemer Reporting

Time: 1500-17:00

Crew Members: Heidi Beemer, Donna Viola

Site Location:  0519731   4251449  and 0520249  4251141

Transit Mode: ATV

Objectives: Recon and collect RMR data for two shelter sites. Collect endolyths along the way.

Results: Two sites listed above where found. Physical appearance data and rock samples were collected from both locations. Nonconformities were measured at first site. We still need to return tomorrow to take photos.

A possible soil collection site was also found for LAMBDA

Lessons Learned: Charge your camera before you leave.


Journalist Report

Sukrit Ranjan Reporting

Today was a cold, cold day. We pushed out EVA after EVA to finish up our projects. The EVA planner project finished today, which is awesome! Endolith and geological sample collecting was also done. After a warm, filling dinner of mountain chilli, the astronomers are getting ready to head out into the bitter night. Ad astra!


Cook's Report

Donna Viola Reporting

02/09/2011

1a) Today was

-A non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Alpine Aire Mountain Chili

Mashed potatoes

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)  (Example:  2 servings potato soup stored; 2 servings rice burned/inedible, discarded)

None

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating. followed directions on package; just add water for both

5)  Comments/questions for Kim and Jean, the food study investigators